# Quasi-Monte Carlo with one categorical variable

### v1

@@ 3.2 Suboptimality
In ansatz 0 we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
@@
The minimax approach of equations (16) gives us ways to choose $\gamma$ for a given list of values $\alpha_{1},\dots,\alpha_{L}$. It often happens that the worst value of $\rho$ for given $\gamma$ is one of the allowed extremes, but this does not always happen. We have seen such exceptions when one of the $\alpha_{\ell}$ is much larger than the others. In our numerical examples the maximum over $\rho$ of $I_{0}(2\!\mid\!\rho)$ or $I_{0}(3\!\mid\!\rho)$ was not much larger than $I_{0}(\gamma_{0}\!\mid\!\rho)$.
Next we consider recommendations that can be made generally without specific values of $\alpha_{\ell}$ in mind. The first thing to notice is that both $I_{0}$ and $I_{1}$ contain a numerator sum that could potentially have a negative power of $\alpha_{\ell}$. For general use where some $\alpha_{\ell}$ might be small, we want to avoid that possibility. There is no possibility of an even more negative power in the denominator to compensate. To keep a positive exponent in the numerator of $I_{0}$ we must have $2-2\rho/(\gamma+1)\geqslant 0$. After rearrangement this becomes a constraint that $\gamma\geqslant\rho-1$, whatever the unknown $\rho$ might be. When $\rho\doteq 3$ is a possibility then we should take $\gamma\geqslant 2$. If instead we use $\gamma=1$ in a $\rho=3$ setting then the first factor in $I_{0}$ is greater than $\alpha_{L}^{-1}$. For $I_{1}$ we want $1-\rho/(\gamma+2)\geqslant 0$ or $\gamma\geqslant\rho-2$. Since we already had $\gamma\geqslant\rho-1$ using ansatz 0, this is not a further constraint. When we know that $\rho\leqslant 2$, then any $\gamma\geqslant 1$ avoids a negative exponent for $\alpha_{\ell}$.
@@ 4 Making each $n_{\ell}$ a power of two
Here we consider how to allocate sample sizes $n_{\ell}=2^{m_{\ell}}$ that sum to given a total sample size $n=2^{m}$. We assume throughout that $n\geqslant L$ and that we want every $n_{\ell}\geqslant 1$. The advantage of having every $n_{\ell}$ be a power of two is that we then get scrambled net sampling within each of the $L$ strata as shown by Proposition 2. Sample sizes that are a power of two are a best practice in scrambled Sobol’ sampling (Owen, 2020). The results with MSE $O(n^{-3+\epsilon})$ for scrambled Sobol’ points require such sample sizes.
The potential disadvantage of having $n_{\ell}$ be a power of $2$ is that we have many fewer sampling ratios $\alpha_{\ell}/\beta_{\ell}$ at our disposal. For example, with $L=2$, we would have to use $n_{1}=n_{2}=n/2$ and this could be a bad fit for $\alpha_{1}=0.9999$ and $\alpha_{2}=0.0001$.
@@
We suppose that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$ for $L\geqslant 2$ and look at sample fractions $(n_{1}/n,n_{2}/n,\dots,n_{L}/n)$. For $L=2$ the only solution is $(1/2,1/2)$ and for $L=3$ the only solution is $(1/2,1/4,1/4)$. For $L=4$, there are two possibilities: $(1/2,1/4,1/8,1/8)$ and $(1/4,1/4,1/4,1/4)$ and for $5\leqslant L\leqslant 10$ the numbers of possibilities are: 3, 5, 9, 16, and 28. The tables in Figure 1 show the partitions for $4\leqslant L\leqslant 8$.
@@
The number of such partitions as a function of $L$ is given by sequence A002572 at the online encyclopedia of integer sequences, (OEIS Foundation Inc., 2023) along with motivating applications such as counting the number of rooted binary trees or, more recently, counting nonequivalent compact Huffman codes (Elsholtz et al. (2023)). The latter reference mentions that the number of solutions is asymptotic to $R\theta^{L}$ for $R\doteq 0.14$ and $\theta\doteq 1.794$. This is a somewhat faster growth rate than the Fibonacci numbers have: their counterpart to $\theta$ is $(1+\sqrt{5})/2\doteq 1.618$. That asymptote is not a good description for small $L$. For $L=10$, it is over 4000, while the exact value is 28.
To choose $\beta_{\ell}$, or equivalently $n_{\ell}$, we uses a forward stepwise algorithm. We initialize every $m_{\ell}=0$ so every $n_{\ell}=2^{m_{\ell}}$ is initially $1$. At every step of the algorithm there are $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$ samples left to allocate. If $n^{\prime}=0$, the algorithm terminates. If $n^{\prime}>0$ then we select one of the indices $\ell$ with $n_{\ell}\leqslant n^{\prime}$ and double that $n_{\ell}$, replacing $m_{\ell}$ by $m_{\ell}+1$. Before describing which $n_{\ell}$ to double, we show that any such algorithm always terminates.
@@
First, if $n^{\prime}>0$ then there is an ‘eligible’ index $\ell$ with $n_{\ell}\leqslant n^{\prime}$, that can be doubled without causing $\sum_{\ell=1}^{L}n_{\ell}>n$. At any stage in the algorithm, $n_{\ell^{*}}$ for $\ell^{*}=\operatorname*{argmin}_{1\leqslant\ell\leqslant L}n_{\ell}$ is a power of two. It divides $n$ and it divides every $n_{\ell}$, so it divides the nonzero value $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$, hence $n_{\ell^{*}}\leqslant n^{\prime}$. Finally, every step with $n^{\prime}>0$ reduces $n^{\prime}$, and so $n^{\prime}$ must reach zero in a finite number of steps starting from $n^{\prime}=n-L$. ∎
@@ 5 Minimax choice
We consider the increase in variance that we get using sample sizes $\bm{n}$ compared to some better sample sizes $\tilde{\bm{n}}$ that we might otherwise have used. For ansatzes $a=0,1$ we define suboptimality ratios
$R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}}{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}\tilde{n}_{\ell}^{-\rho/2}}\quad\text{and}\quad R_{0}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}n_{\ell}^{-\rho}}{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}\tilde{n}_{\ell}^{-\rho}}$
@@ 6 A toy example
The within-stratum integrand is $g(x_{1})=\exp(-x_{1}^{2})\cos(x_{1})$. The integrand $f(\bm{z})$ for $\bm{z}\in[0,1]^{2}$ has a first component $z_{1}$ that selects a stratum $\ell$ as described in Section 2.2. Then $x_{1}=\theta_{\ell}+\Phi^{-1}(z_{2})$ where $\Phi(\cdot)$ is the cumulative distribution function of the $\mathcal{N}(0,1)$ distribution. We chose this function because it is regular enough that we expect to see an RQMC variance of $O(n^{-3})$.
@@
    An importance adjusted estimator: $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})\omega(\ell(i))$ with RQMC points where the $n_{\ell}$ come from rounding equation (14), using $\rho=2$, because that is the rate from the K-H bound for $n_{\ell}$ not equal to powers of two.
@@
For each estimator we compute $R=500$ replicates at sample sizes $n=2^{m}$ for $3\leqslant m\leqslant 12$ and then graph the sample variances.
@@
The variance for Monte Carlo sampling follows the expected $O(1/n)$ rate. Simply replacing the MC points by RQMC points improves the rate to roughly $O(1/n^{2})$. Adjusting the sampling fractions from $\alpha_{\ell}$ to $\beta_{\ell}$ using $\rho=2$ retains the RQMC rate and provides lower variance at larger samples sizes. Using $\rho=3$ to choose sample sizes $n_{\ell}$ we compared two methods. The estimate $\hat{\mu}_{\mathrm{RQMC}(L)}$ uses $L$ independent RQMC samples, with $n_{\ell}$ points in stratum $\ell$. The other uses the combined RQMC estimate $\hat{\mu}_{\mathrm{RQMC}(2)}$. These latter two both give an apparent $O(n^{-3})$ rate for the variance. The combined estimate has a better constant in this example. See our discussion of between strata correlations in Section 8.
@@ 7 Saint-Venant flood model
where $Q$ is the flow rate (meters3/second), $B$ is the width of the river (meters), $K_{s}$ is the Strickler coefficient (meters1/3/second) quantifying the roughness of the river bed and $\alpha$ is the (dimensionless) slope of the river bed. It is given by $\alpha=(Z_{m}-Z_{v})/L$ where $Z_{m}$ and $Z_{v}$ are the upstream and downstream heights of the river in meters.
@@
This model is usually given with a symmetric triangular density for $L$ over the interval $[4990,5010]$. It is usually seen that uncertainty in $L$ makes very little difference. See for example Iooss and Lemaître (2015). Therefore we fix $L=5000$. Similarly, $B$ is commonly given a symmetric triangular distribution over $[295,305]$ and we fix it at $B=300$. We consider four distributions for the inputs $Q$, $K_{s}$, $Z_{v}$ and $Z_{m}$. With probability $0.95$ all four random variables are independently distributed with nominal distributions given in Table 2. With probability $0.02$, the distribution of $Q$ takes the adverse distribution in Table 2 while the other variables are nominal. With probability $0.02$, the distribution of $K_{s}$ is adverse and the others are nominal. With probability $0.01$ both $Q$ and $K_{s}$ are adverse.
The distribution of $K_{s}$ is customarily given as $\mathcal{N}(30,7.5^{2})$ and then truncated to a range such as $(0,\infty)$ in Limbourg and De Rocquigny (2010) or $(15,\infty)$ in Iooss (2011). We have chosen a nominal Gamma distribution, so we don’t need any truncation. That Gamma distribution has mean $30$ but a lower variance than $7.5^{2}$. The adverse distribution for the Gamma distribution takes lower values which bring more flooding as $K_{s}$ appears in the denominator of $H$. The distribution of $Q$ is customarily given as Gumbel with mode $1013$ and scale $558$. This was found by a maximum likelihood fit, reported in Limbourg and De Rocquigny (2010). The Gumbel distribution has an extreme value justification, but it also gives negative values which are unphysical. We have replaced it by a Fréchet distribution which is also an extreme value distribution. It has much heavier tails than the Gumbel which one might need to characterize extreme risks. Our nominal Fréchet distribution has a similar mean to the Gumbel distribution above. The adverse distribution triples the nominal one.
@@
In Figure 4 we consider integer allocations with different values of $\rho$ using equation 14 to select $n_{\ell}$. For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope $-1.8$. The curve for $\rho=2$ which is also in Figure 3 is best at larger $n$. The case of $\rho=1$ has proportional allocation and may have an empty stratum for $n\leqslant 128$ (as $128*0.01<2$).

### v2

@@ 3.2 Suboptimality
In ansatz 0, we would then take $n_{\ell}\propto\alpha_{\ell}^{2/(\gamma+1)}$. The inefficiency of using $\gamma$ given that we should have used $\rho$ is
@@
The minimax approach of equation (16) gives us ways to choose $\gamma$ for a given list of values $\alpha_{1},\dots,\alpha_{L}$. It often happens that the worst value of $\rho$ for given $\gamma$ is at one of the allowed extremes, but this does not always happen. We have seen such exceptions when one of the $\alpha_{\ell}$ is much larger than the others. In our numerical examples the maximum over $\rho$ of $I_{0}(2\!\mid\!\rho)$ or $I_{0}(3\!\mid\!\rho)$ was not much larger than $I_{0}(\gamma_{0}\!\mid\!\rho)$.
Next we consider recommendations that can be made generally without specific values of $\alpha_{\ell}$ in mind. The first thing to notice is that both $I_{0}$ and $I_{1}$ contain a numerator sum that could potentially have a negative power of $\alpha_{\ell}$. For general use where some $\alpha_{\ell}$ might be small, we want to avoid that possibility. There is no possibility of an even more negative power in the denominator to compensate. To keep a positive exponent in the numerator of $I_{0}$ we must have $2-2\rho/(\gamma+1)\geqslant 0$. After rearrangement this becomes a constraint that $\gamma\geqslant\rho-1$, whatever the unknown $\rho$ might be. When $\rho\doteq 3$ is a possibility, then we should take $\gamma\geqslant 2$. If instead we use $\gamma=1$ in a $\rho=3$ setting then the first factor in $I_{0}$ is greater than $\alpha_{L}^{-1}$. For $I_{1}$ we want $1-\rho/(\gamma+2)\geqslant 0$ or $\gamma\geqslant\rho-2$. Since we already had $\gamma\geqslant\rho-1$ using ansatz 0, this is not a further constraint. When we know that $\rho\leqslant 2$, then any $\gamma\geqslant 1$ avoids a negative exponent for $\alpha_{\ell}$.
@@ 4 Making each $n_{\ell}$ a power of two
Here we consider how to allocate sample sizes $n_{\ell}=2^{m_{\ell}}$ that sum to a given total sample size $n=2^{m}$. We assume throughout that $n\geqslant L$ and that we want every $n_{\ell}\geqslant 1$. The advantage of having every $n_{\ell}$ be a power of two is that we then get scrambled net sampling within each of the $L$ strata as shown by Proposition 2. Sample sizes that are a power of two are a best practice in scrambled Sobol’ sampling (Owen, 2020). The results with MSE $O(n^{-3+\epsilon})$ for scrambled Sobol’ points require such sample sizes.
A potential non-asymptotic disadvantage of having $n_{\ell}$ be a power of $2$ is that we have many fewer sampling ratios $\alpha_{\ell}/\beta_{\ell}$ at our disposal. For example, with $L=2$, we would have to use $n_{1}=n_{2}=n/2$ and this could be a bad fit for $\alpha_{1}=0.9999$ and $\alpha_{2}=0.0001$.
@@
We suppose that $\beta_{1}\geqslant\beta_{2}\geqslant\cdots\geqslant\beta_{L}$ for $L\geqslant 2$ and look at sample fractions $(n_{1}/n,n_{2}/n,\dots,n_{L}/n)$. For $L=2$ the only solution is $(1/2,1/2)$ and for $L=3$ the only solution is $(1/2,1/4,1/4)$. For $L=4$, there are two possibilities: $(1/2,1/4,1/8,1/8)$ and $(1/4,1/4,1/4,1/4)$ and for $5\leqslant L\leqslant 10$ the numbers of possibilities are: 3, 5, 9, 16, 28 and 50. The tables in Figure 1 show the partitions for $4\leqslant L\leqslant 8$.
@@
The number of such partitions as a function of $L$ is given by sequence A002572 at the online encyclopedia of integer sequences, (OEIS Foundation Inc., 2023) along with motivating applications such as counting the number of rooted binary trees or, more recently, counting nonequivalent compact Huffman codes (Elsholtz et al. (2023)). The latter reference mentions that the number of solutions is asymptotic to $R\theta^{L}$ for $R\doteq 0.14$ and $\theta\doteq 1.794$. This is a somewhat faster growth rate than the Fibonacci numbers have: their counterpart to $\theta$ is $(1+\sqrt{5})/2\doteq 1.618$. For $L=10$, the approximation is close to 48.3.
To choose $\beta_{\ell}$, or equivalently $n_{\ell}$, we use a forward stepwise algorithm. We initialize every $m_{\ell}=0$ so every $n_{\ell}=2^{m_{\ell}}$ is initially $1$. At every step of the algorithm there are $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$ samples left to allocate. If $n^{\prime}=0$, the algorithm terminates. If $n^{\prime}>0$ then we select one of the indices $\ell$ with $n_{\ell}\leqslant n^{\prime}$ and double that $n_{\ell}$, replacing $m_{\ell}$ by $m_{\ell}+1$. Before describing which $n_{\ell}$ to double, we show that any such algorithm always terminates.
@@
First, if $n^{\prime}>0$ then we can show that there is an ‘eligible’ index $\ell$ with $n_{\ell}\leqslant n^{\prime}$, that can be doubled without causing $\sum_{\ell=1}^{L}n_{\ell}>n$. At any stage in the algorithm, $n_{\ell^{*}}$ for $\ell^{*}=\operatorname*{argmin}_{1\leqslant\ell\leqslant L}n_{\ell}$ is a power of two. It divides $n$ and it divides every $n_{\ell}$, so it also divides the nonzero value $n^{\prime}=n-\sum_{\ell=1}^{L}n_{\ell}$, hence $n_{\ell^{*}}\leqslant n^{\prime}$. Finally, every step with $n^{\prime}>0$ reduces $n^{\prime}$, and so $n^{\prime}$ must reach zero in a finite number of steps starting from $n^{\prime}=n-L$. ∎
@@ 5 Minimax choice
We consider the increase in variance that we get using sample sizes $\bm{n}$ compared to some better sample sizes $\tilde{\bm{n}}$ that we might otherwise have used. For ansatzes 00 and $1$ we define suboptimality ratios
$R_{0}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}n_{\ell}^{-\rho}}{\sum_{\ell}\alpha_{\ell}^{2}\tau_{\ell}\tilde{n}_{\ell}^{-\rho}}\quad\text{and}\quad R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau}):=\frac{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}n_{\ell}^{-\rho/2}}{\sum_{\ell}\alpha_{\ell}\tau_{\ell}^{1/2}\tilde{n}_{\ell}^{-\rho/2}}$
@@ 6 A toy example
The within-stratum integrand is $g(x_{1})=\exp(-x_{1}^{2})\cos(x_{1})$. The integrand $f(\bm{z})$ for $\bm{z}\in[0,1]^{2}$ has a first component $z_{1}$ that selects a stratum $\ell$ as described in Section 2.2. Then $x_{1}=\theta_{\ell}+\Phi^{-1}(z_{2})$ where $\Phi(\cdot)$ is the cumulative distribution function of the $\mathcal{N}(0,1)$ distribution. We chose this function because it is regular enough that we expect to see an RQMC variance of $O(n^{-3})$. We chose these values of $\alpha_{\ell}$ so that the forward algorithm would give different $n_{\ell}$ for $\rho=2$ than it does for $\rho=3$.
@@
    An importance adjusted estimator: $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})\omega(\ell(i))$ with RQMC points where the $n_{\ell}$ come from rounding equation (14), using $\rho=2$, because that is the rate from the Koksma-Hlawka bound for $n_{\ell}$ not equal to powers of two.
@@
For each estimator we compute $R=500$ replicates at sample sizes $n=2^{m}$ for $3\leqslant m\leqslant 12$ and then graph the sample variances. We used nested uniform scrambling (Owen, 1995) to do the scrambling.
@@
The variance for Monte Carlo sampling follows the expected $O(1/n)$ rate. Simply replacing the MC points by RQMC points improves the rate to roughly $O(1/n^{2})$. Adjusting the sampling fractions from $\alpha_{\ell}$ to $\beta_{\ell}$ using $\rho=2$ retains the RQMC rate and provides lower variance at larger samples sizes. Using $\rho=3$ to choose sample sizes $n_{\ell}$, we compared two methods. The first generates the estimate $\hat{\mu}_{\mathrm{RQMC}(L)}$ using $L$ independent RQMC samples, with $n_{\ell}$ points in stratum $\ell$. The other uses the combined RQMC estimate $\hat{\mu}_{\mathrm{RQMC}(2)}$. These latter two both give an apparent $O(n^{-3})$ rate for the variance. The combined estimate has a better constant in this example. See our discussion of between strata correlations in Section 8.
@@ 7 Saint-Venant flood model
where $Q$ is the flow rate (meters3/second), $B$ is the width of the river (meters), $K_{s}$ is the Strickler coefficient (meters1/3/second) quantifying the roughness of the river bed and $\alpha$ is the (dimensionless) slope of the river bed. It is given by $\alpha=(Z_{m}-Z_{v})/D$ where $Z_{m}$ and $Z_{v}$ are the upstream and downstream heights of the river in meters, and $D$ denotes the river’s length in meters.
@@
This model is usually given with a symmetric triangular density for $D$ over the interval $[4990,5010]$. It is usually seen that uncertainty in $D$ makes very little difference. See for example Iooss and Lemaître (2015). Therefore we fix $D=5000$. Similarly, $B$ is commonly given a symmetric triangular distribution over $[295,305]$ and we fix it at $B=300$. We consider four distributions for the inputs $Q$, $K_{s}$, $Z_{v}$ and $Z_{m}$. With probability $0.95$ all four random variables are independently distributed with nominal distributions given in Table 2. With probability $0.02$, the distribution of $Q$ takes the adverse distribution in Table 2 while the other variables are nominal. With probability $0.02$, the distribution of $K_{s}$ is adverse and the others are nominal. With probability $0.01$ both $Q$ and $K_{s}$ are adverse.
The distribution of $K_{s}$ is customarily given as $\mathcal{N}(30,7.5^{2})$ and then truncated to a range such as $(0,\infty)$ in Limbourg and De Rocquigny (2010) or $(15,\infty)$ in Iooss (2011). We have chosen a nominal Gamma distribution, so we do not need any truncation. That Gamma distribution has mean $30$ but a lower variance than $7.5^{2}$. The adverse distribution for the Gamma distribution takes lower values which bring more flooding as $K_{s}$ appears in the denominator of $H$. The distribution of $Q$ is customarily given as Gumbel with mode $1013$ and scale $558$. This was found by a maximum likelihood fit, reported in Limbourg and De Rocquigny (2010). The Gumbel distribution has an extreme value justification, but it also gives negative values which are unphysical. We have replaced it by a Fréchet distribution which is always positive and is also an extreme value distribution. It has much heavier tails than the Gumbel which one might need to characterize extreme risks. Our nominal Fréchet distribution has a similar mean to the Gumbel distribution above. The adverse distribution triples the nominal one.
@@
In Figure 4 we consider integer allocations with different values of $\rho$ using equation (14) to select $n_{\ell}$. For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope close to $-1.8$. The curve for $\rho=2$ which is also in Figure 3 is best at larger $n$. The case of $\rho=1$ has proportional allocation and may have an empty stratum for $n\leqslant 128$ (as $128*0.01<2$).
@@ 8 Discussion
We might also have a variable with a mixed distribution including a continuous component with a density function and $L$ atoms. For this setting we can make the continuous part of that variable’s distribution into an $L+1$’st mixture component. We will still need to sample within the range of that continuous component. For that we can add an $s+1$’st component to $\bm{u}$, which is then the $s+2$’nd component of $\bm{z}$.